OpenStudy (anonymous):

Select the coordinates of two points on the line y = 8. A.) (8, 8) and (8, 0) B.)(8, -8) and (8, 8) C.)(0, 8) and (8, 0) D.) (0, 8) and (8, 8)

OpenStudy (imstuck):

If the line is a horizontal line, then all the y points will be the same along the line. They will all be an 8. Which two sets of coordinates you gave had an 8 in the y position? Not a negative 8, either. They are both 8.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Horizontal line...y coordinate must be 8

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

I have some other questions if you can help me.

OpenStudy (anonymous):

Graph with a line going through points negative 2 comma zero and 0 comma negative one. Select the equation of a line that is perpendicular to the line on the graph and passes through the point (-1, 2). y = 2x + 4 y = - 2x + 2 y = - 1 over 2 x + 2 y = 2x - 1

OpenStudy (anonymous):

The slope of the line containing the points (-2,0)(0,-1) is -1/2 work: 0--1/-2-0 = 1/-2 or -1/2 To have a line perpendicular, the slope is the opposite inverse of the given line. So the slope of the perpendicular line is 2. Point (-1,2) Use the equation y - y1 = m(x-x1) y - 2 = 2(x--1) y - 2 = 2x + 2 y = 2x + 4 This is the line perpendicular to the given line that passes through the point (-1,2)

OpenStudy (anonymous):

So the answer will be C

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Let me look at this again

OpenStudy (anonymous):

It's A...y = 2x + 4

OpenStudy (anonymous):

Thank you for your help.